# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing, but their susceptibility to biases poses significant challenges. This comprehensive review examines the landscape of bias in LLMs, from its origins to current mitigation strategies. We categorize biases as intrinsic and extrinsic, analyzing their manifestations in various NLP tasks. The review critically assesses a range of bias evaluation methods, including data-level, model-level, and output-level approaches, providing researchers with a robust toolkit for bias detection. We further explore mitigation strategies, categorizing them into pre-model, intra-model, and post-model techniques, highlighting their effectiveness and limitations. Ethical and legal implications of biased LLMs are discussed, emphasizing potential harms in real-world applications such as healthcare and criminal justice. By synthesizing current knowledge on bias in LLMs, this review contributes to the ongoing effort to develop fair and responsible AI systems. Our work serves as a comprehensive resource for researchers and practitioners working towards understanding, evaluating, and mitigating bias in LLMs, fostering the development of more equitable AI technologies.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs) has transformed the field of Natural Language Processing (NLP), introducing new possibilities and applications across various domains, including healthcare (Keskar et al., 2019), finance (Yang et al., 2020), education (Okonkwo and Ade-Ibijola, 2023), and entertainment (Jiang et al., 2020). These models, such as GPT-3 (Brown et al., 2020), BERT (Devlin et al., 2018), and others, are designed to understand and generate human-like text by learning from vast corpora of text data. These models have become the backbone for tasks such as machine translation (Lewis et al., 2020), text summarization (Nallapati et al., 2016), sentiment analysis (Zhang et al., 2018), and automated question-answering (Raffel et al., 2020), reflecting their deep integration into various industries and day-to-day applications, and making them indispensable tools in the modern AI landscape.
@@
Despite their immense potential and utility, LLMs have raised concerns due to their inherent biases that reflect societal prejudices present in their training data (Bender et al., 2021; Blodgett et al., 2020). These biases, which manifest as gender, racial, cultural, and socio-economic stereotypes, pose serious ethical and practical challenges, especially when LLMs are deployed in critical decision-making environments such as healthcare diagnostics (Rajkomar et al., 2018), legal judgments (Angwin et al., 2016), and hiring processes (Chen et al., 2018). Such biases can lead to unequal treatment or skewed results that disproportionately affect marginalized groups, potentially exacerbating existing inequalities. Caliskan et al. (2017) has noted that LLMs often mirror historical biases present in human language, risking the amplification of these biases within automated systems.
@@
Understanding bias within LLMs requires insight into how biases traditionally arise in statistical models. In classical statistics, biases are well-studied in models for continuous, binary, and time-to-event endpoints Hastie et al. (2009). These biases manifest due to several issues, such as selection bias Akaike (1974); Heckman (1979); Vardi (1982), overfitting Heckman (1979), handling of covariates Akaike (1974); Tibshirani (1996) , limited sample sizes Nemes et al. (2009), and even randomization ratio Hongfei Li and Hou (2024a, b), which can all lead to incorrect estimates and predictions. These statistical biases share important parallels with those encountered in LLMs. For instance, LLMs must handle incomplete or unbalanced datasets carefully to prevent skewed outputs. Similarly, the challenges of selection bias in classical models echo the biases found in LLMs trained on non-representative data sources, which can lead to harmful disparities in model predictions, especially for underrepresented groups. Addressing these statistical biases requires a range of mitigation strategies, including robust causal inference framework Rubin (1974). Similarly, bias mitigation in LLMs involves data-level interventions such as resampling and augmentation, model-level adjustments with fairness constraints, and post-processing corrections to fine-tune outputs (Schick et al., 2021).
@@
LLMs also exhibit biases similar to those observed in traditional statistical models. Bias in LLMs can be broadly categorized into two types: Intrinsic bias and Extrinsic bias. Intrinsic biases originate from the training data, as well as the architecture and underlying assumptions made during model design (Sun et al., 2019). LLMs, trained on vast datasets often sourced from internet and textual repositories, inevitably inherit biases present in these sources. For example, language models may reinforce gender stereotypes in occupations, associating “doctor” with men and “nurse” with women (Bolukbasi et al., 2016; Zhao et al., 2018a). Similarly, certain demographic groups may be either underrepresented or misrepresented within training datasets, further compounding model biases (Hovy and Prabhumoye, 2021). This becomes particularly problematic when LLMs are applied in sensitive domains where accurate and unbiased information is crucial for decision-making (Bender et al., 2021).
@@
Extrinsic Biases, on the other hand, emerge during the application of LLMs in real-world tasks. These biases are often more subtle as they manifest in the model outputs during specific tasks, such as sentiment analysis, content moderation, or automated decision-making systems. For example, Sap et al. (2019a) found that LLMs used in hate speech detection could disproportionately flag certain dialects or vernaculars, such as African American Vernacular English (AAVE), as more offensive compared to standardized English, reinforcing societal stereotypes. Similarly, Kiritchenko and Mohammad (2018) demonstrated that sentiment analysis systems could produce biased results when analyzing texts associated with different demographic groups, thereby perpetuating harmful stereotypes in their predictions.
@@
The implications of biases in LLMs are far-reaching. In healthcare, for instance, biased models can lead to inappropriate treatment recommendations, potentially worsening existing health inequalities between demographic groups (Obermeyer et al., 2019). In legal settings, reliance on biased language models for risk assessment or sentencing decisions can result in discriminatory practices (Angwin et al., 2016). Moreover, biases embedded in LLMs affect everyday applications, such as search engines and social media platforms, where they shape public discourse, potentially reinforcing echo chambers and marginalizing minority voices (Pariser, 2011; Sap et al., 2019b).
@@
These biases not only perpetuate inequality but also raise significant ethical and legal questions regarding the use of LLMs in decision-making processes. Given the critical nature of these challenges, there is an urgent need for comprehensive evaluation frameworks to detect, quantify, and mitigate these biases, ensuring that LLMs function as fair and equitable tools for all users. Sheng et al. (2021) provide an overview of these efforts, noting that bias evaluation typically involves analyzing models at various stages of their lifecycle—from data preprocessing to model training and output generation. Common techniques include the use of benchmark datasets designed to highlight biases (Nadeem et al., 2021), probing models to understand their internal representations (Hewitt and Manning, 2019), and conducting user studies to assess the real-world impact of biased outputs (Sap et al., 2019b).
@@
In terms of bias mitigation, strategies are typically categorized into three broad approaches: data-level interventions, model-level techniques, and post-processing adjustments. Data-level interventions involve curating and balancing the training datasets to reduce the prevalence of biased content, often through data augmentation, filtering, or resampling techniques (Zhao et al., 2018b). However, this approach is limited by the vast scale and inherent diversity of LLM training datasets, making it difficult to eliminate biases entirely. Model-level techniques may include altering the training objectives, incorporating fairness constraints, or modifying model architectures to minimize biased learning (Liang et al., 2020b). These methods often trade off some degree of model accuracy for fairness, presenting challenges in real-world deployment. Post-processing methods focus on adjusting the outputs of LLMs after generation, applying debiasing algorithms, controlled text generation, or reinforcement learning frameworks to reduce biased content (Schick et al., 2021). While this approach offers flexibility, it may require substantial computational overhead and may not fully address the root causes of bias.
@@
Despite these advancements, fully addressing bias in LLMs remains a complex and ongoing challenge. The complexity of language, the vastness of training datasets, and the massive architectures of modern LLMs–often comprising billions of parameters–make it difficult to identify and correct all instances of bias. Furthermore, as Mehrabi et al. (2021) suggest, biases are often context-dependent and may evolve over time as societal norms and values change. This dynamic nature of bias necessitates continuous monitoring, evaluation, and refinement of both models and the methodologies used to assess their outputs.
@@
This article aims to provide a comprehensive exploration of bias in LLMs, examining its origins, manifestations, and the current state of research in bias evaluation and mitigation. We begin by discussing the sources and types of biases, drawing on examples from various NLP tasks to illustrate how biases arise and persist. We then review the methodologies used to detect and measure biases, including both quantitative and qualitative approaches. Following this, we explore the strategies developed to mitigate biases at different stages of the model development pipeline, assessing their effectiveness and limitations. Finally, we discuss the ethical considerations and legal implications of deploying biased AI systems, underscoring the importance of fairness, transparency, and accountability in AI development. Through this discussion, we seek to contribute to the broader discourse on responsible AI, offering insights and recommendations for researchers, practitioners, and policymakers working to ensure that LLMs serve as equitable and trustworthy tools in society.
@@ 2 Intrinsic Bias
The bias presented in Large Language Models (LLMs) can be broadly categorized into intrinsic bias and extrinsic bias based on the different stages at which the biases manifest within the model’s lifecycle and the type of bias being measured (Doan et al., 2024). Intrinsic bias refers to biases that are inherently within the internal representations or outputs of a trained or pre-trained LLM and are independent of any specific downstream tasks. Extrinsic bias refers to the biases that manifest during the model’s performance on specific downstream tasks after training or fine-turing. Intrinsic biases are typically encoded during the training or pre-training phase when the LLM learns from large-scale corpora. For instance, if a pre-trained LLM consistently associates certain professions with specific genders (e.g., assuming all doctors to be male and assuming all nurses to be female) or racial stereotypes (May et al., 2019), this bias would be considered intrinsic bias. This type of bias is tied to the way the model has encoded relationships and patterns in its internal representations. In this section, we introduce the possible sources of intrinsic bias, present the manifestations of intrinsic bias, and show available measurement and detection of Intrinsic bias.
2.1 Sources of Intrinsic Bias
@@
Intrinsic biases are primarily introduced during the training/pre-training stages where models learn patterns and representations from vast corpora. These biases can be traced to several aspects of data and model design, which will be outlined below.
2.1.1 Bias in training Data
@@
Language models, especially LLMs that are trained on large corpus, usually have intrinsic bias issues, since the training corpora often contains societal biases that are built into the model (Pagano et al., 2023; Ray, 2023; Goldfarb-Tarrant, 2024; Goldman and Tsotsos, 2024). These biases are present in several ways shown as follows.
@@
  • Bias from over-representativeness and under-representativeness: Certain groups, such as gender, age, race, religion, ethnicity, culture, political, or socioeconomic classes, may be under-represented or over-represented in the corpus (Das et al., 2024; Alvero et al., 2024). For example, men might be over-represented in datasets about leadership or science, while women may be more frequently mentioned in caregiving roles (into Bias, 2024), thus leading to biased associations with demographic information. Another example is that the Latinx population, especially Mexican-American students, remains underrepresented in U.S. higher education (Alvero et al., 2024).
@@
  • Spatial and Temporal Bias: LLMs trained predominantly on the corpus from certain countries or geographic locations may absorb the cultural norms and values, hence building biases into the underlying LLMs. For instance, a model trained on Western-centric data may have a skewed understanding of non-Western cultures (Ahmad and Bhattacharyya, ; Talat et al., 2022; Mukherjee et al., 2024; Lee et al., 2024; Shi et al., 2024). Similarly, data collected over different periods may reflect outdated societal norms and values. LLMs trained on these data may be biased by those outdated norms and values. For instance, historical texts may exhibit racist or sexist terminology, which the model may absorb as part of its internal representations (Salinas et al., 2023; Bussaja, 2024).
2.1.2 Bias in Data Collection Methods

### v2

@@ Abstract
Large language models (LLMs) have revolutionized natural language processing, but their susceptibility to biases poses significant challenges. This comprehensive review examines the landscape of bias in LLMs, from its origins to current mitigation strategies. We categorize biases as intrinsic and extrinsic, analyzing their manifestations in various natural language processing (NLP) tasks. The review critically assesses a range of bias evaluation methods, including data-level, model-level, and output-level approaches, providing researchers with a robust toolkit for bias detection. We further explore mitigation strategies, categorizing them into pre-model, intra-model, and post-model techniques, highlighting their effectiveness and limitations. Ethical and legal implications of biased LLMs are discussed, emphasizing potential harms in real-world applications such as healthcare and criminal justice. By synthesizing current knowledge on bias in LLMs, this review contributes to the ongoing effort to develop fair and responsible artificial intelligence (AI) systems. Our work serves as a comprehensive resource for researchers and practitioners working towards understanding, evaluating, and mitigating bias in LLMs, fostering the development of more equitable AI technologies.
Keywords: bias evaluation; bias mitigation strategies; large language models
@@ 1 Introduction
The rapid development of large language models (LLMs) has transformed the field of natural language processing (NLP), introducing new possibilities and applications across various domains, including healthcare (Keskar et al., 2019), finance (Yang et al., 2020), education (Okonkwo and Ade-Ibijola, 2023), and entertainment (Jiang et al., 2020). These models, such as GPT-3 (Brown et al., 2020), BERT (Devlin et al., 2018), and others, are designed to understand and generate human-like text by learning from vast corpora of text data. These models have become the backbone for tasks such as machine translation (Lewis et al., 2020), text summarization (Nallapati et al., 2016), sentiment analysis (Zhang et al., 2018), and automated question answering (Raffel et al., 2020), reflecting their deep integration into various industries and day-to-day applications, and making them indispensable tools in the modern AI landscape.
@@
Despite their immense potential and utility, LLMs have raised concerns due to their inherent biases that reflect societal prejudices present in their training data (Bender et al., 2021; Blodgett et al., 2020). These biases, which manifest as gender, racial, cultural, and socioeconomic stereotypes, pose serious ethical and practical challenges, especially when LLMs are deployed in critical decision-making environments such as healthcare diagnostics (Rajkomar et al., 2018), legal judgments (Angwin et al., 2016), and hiring processes (Chen et al., 2018). Such biases can lead to unequal treatment or skewed results that disproportionately affect marginalized groups, potentially exacerbating existing inequalities. Caliskan et al. (2017) have noted that LLMs often mirror historical biases present in human language, risking the amplification of these biases within automated systems.
@@
Understanding bias within LLMs requires insight into how biases traditionally arise in statistical models. In classical statistics, biases have been well studied in models for continuous, binary, and time-to-event endpoints (Hastie et al., 2009). These biases manifest due to several issues, such as selection bias (Akaike, 1974; Heckman, 1979; Vardi, 1982), overfitting (Heckman, 1979), handling of covariates (Akaike, 1974; Tibshirani, 1996), limited sample sizes (Nemes et al., 2009), and even the randomization ratio (Hongfei Li and Hou, 2024), which can all lead to incorrect estimates and predictions. These statistical biases share important parallels with those encountered in LLMs. For instance, LLMs must handle incomplete or unbalanced datasets carefully to prevent skewed outputs. Similarly, the challenges of selection bias in classical models echo the biases found in LLMs trained on non-representative data sources, which can lead to harmful disparities in model predictions, especially for under-represented groups. Addressing these statistical biases requires a range of mitigation strategies, including a robust causal inference framework Rubin (1974). Similarly, bias mitigation in LLMs involves data-level interventions such as resampling and augmentation, model-level adjustments with fairness constraints, and post-processing corrections to fine-tune outputs (Schick et al., 2021).
@@
LLMs also exhibit biases similar to those observed in traditional statistical models. Bias in LLMs can be broadly categorized into two types: intrinsic bias and extrinsic bias. Intrinsic biases originate from the training data, as well as the architecture and underlying assumptions made during model design (Sun et al., 2019). LLMs, trained on vast datasets, often sourced from internet and textual repositories, inevitably inherit biases present in these sources. For example, language models may reinforce gender stereotypes in occupations, associating “doctor” with men and “nurse” with women (Bolukbasi et al., 2016; Zhao et al., 2018). Similarly, certain demographic groups may be either under-represented or misrepresented within training datasets, further compounding model biases (Hovy and Prabhumoye, 2021). This becomes particularly problematic when LLMs are applied in sensitive domains where accurate and unbiased information is crucial for decision making (Bender et al., 2021).
@@
Extrinsic biases, on the other hand, emerge during the application of LLMs in real-world tasks. These biases are often more subtle as they manifest in the model outputs during specific tasks, such as sentiment analysis, content moderation, or automated decision-making systems. For example, Sap et al. (2019) found that LLMs used in hate speech detection could disproportionately flag certain dialects or vernaculars, such as African-American Vernacular English (AAVE), as more offensive compared to standardized English, reinforcing societal stereotypes. Similarly, Kiritchenko and Mohammad (2018) demonstrated that sentiment analysis systems could produce biased results when analyzing texts associated with different demographic groups, thereby perpetuating harmful stereotypes in their predictions.
@@
The implications of biases in LLMs are far-reaching. In healthcare, for instance, biased models can lead to inappropriate treatment recommendations, potentially worsening existing health inequalities between demographic groups (Obermeyer et al., 2019). In legal settings, reliance on biased language models for risk assessment or sentencing decisions can result in discriminatory practices (Angwin et al., 2016). Moreover, biases embedded in LLMs affect everyday applications, such as search engines and social media platforms, where they shape public discourse, potentially reinforcing echo chambers and marginalizing minority voices (Pariser, 2011; Sap et al., 2019).
@@
These biases not only perpetuate inequality but also raise significant ethical and legal questions regarding the use of LLMs in decision-making processes. Given the critical nature of these challenges, there is an urgent need for comprehensive evaluation frameworks to detect, quantify, and mitigate these biases, ensuring that LLMs function as fair and equitable tools for all users. Sheng et al. (2021) provide an overview of these efforts, noting that bias evaluation typically involves analyzing models at various stages of their lifecycle—from data preprocessing to model training and output generation. Common techniques include the use of benchmark datasets designed to highlight biases (Nadeem et al., 2021), probing models to understand their internal representations (Hewitt and Manning, 2019), and conducting user studies to assess the real-world impact of biased outputs (Sap et al., 2019).
@@
In terms of bias mitigation, strategies are typically categorized into three broad approaches: data-level interventions, model-level techniques, and post-processing adjustments. Data-level interventions involve curating and balancing the training datasets to reduce the prevalence of biased content, often through data augmentation, filtering, or resampling techniques (Zhao et al., 2018). However, this approach is limited by the vast scale and inherent diversity of LLM training datasets, making it difficult to eliminate biases entirely. Model-level techniques may include altering the training objectives, incorporating fairness constraints, or modifying model architectures to minimize biased learning (Liang et al., 2020). These methods often trade off some degree of model accuracy for fairness, presenting challenges in real-world deployment. Post-processing methods focus on adjusting the outputs of LLMs after generation, applying debiasing algorithms, controlled text generation, or reinforcement learning frameworks to reduce biased content (Schick et al., 2021). While this approach offers flexibility, it may require substantial computational overhead and may not fully address the root causes of bias.
@@
Despite these advancements, fully addressing bias in LLMs remains a complex and ongoing challenge. The complexity of language, the vastness of training datasets, and the massive architectures of modern LLMs—often comprising billions of parameters—make it difficult to identify and correct all instances of bias. Furthermore, as Mehrabi et al. (2021) suggest, biases are often context-dependent and may evolve over time as societal norms and values change. This dynamic nature of bias necessitates continuous monitoring, evaluation, and refinement of both models and the methodologies used to assess their outputs.
To boost the reproducibility of this review and allow readers to audit the review’s coverage, the literature collection process is documented as follows. Specifically, we searched ACL Anthology, arXiv (cs.CL, cs.AI, cs.LG), Google Scholar, Semantic Scholar, and ACM Digital Library, covering major NLP and AI venues, including ACL, EMNLP, NAACL, NeurIPS, ICML, AAAI, and FAccT. Primary query strings included combinations of (“large language model” OR “LLM” OR “language model”) AND (“bias” OR “fairness” OR “stereotype” OR “debiasing”). The search covered publications from January 2016 through August 2024. Inclusion criteria required papers to directly address bias origins, evaluation, or mitigation in language models, be published in peer-reviewed venues or as substantive preprints, and be written in English. Approximately 200 papers were included in the final synthesis after title, abstract, and full-text screening.
This article provides a structured synthesis of current research on bias in LLMs, covering its origins, manifestations, evaluation, mitigation, and broader societal implications. While prior reviews such as Mehrabi et al. Mehrabi et al. (2021), Navigli et al. Navigli et al. (2023), and Gallegos et al. Gallegos et al. (2024) have made important contributions to bias phenomena and mitigation methods, the present review distinguishes itself through five specific contributions. First, we propose and operationalize a precise two-level taxonomic framework distinguishing intrinsic bias, which is encoded within model representations during pretraining, from extrinsic bias, which emerges in downstream tasks and generated outputs, providing clearer conceptual boundaries than existing categorizations. Second, we review bias evaluation methods across data-, model-, and output-level analyses, while also discussing human-involved and domain-specific evaluation perspectives. Third, we provide a comparative synthesis of pre-model, intra-model, and post-model mitigation strategies, explicitly examining trade-offs among computational cost, generalizability, and fairness that have received less attention in prior surveys. Fourth, we examine the ethical and legal implications of biased LLMs through a structured taxonomy of representational and allocational harms, grounding abstract fairness concerns in documented real-world consequences. Fifth, we highlight cross-cultural and multilingual gaps that remain under-represented in the current literature, documenting bias phenomena and evaluation gaps across East Asian, Arabic, South Asian, African, and Latin American linguistic and cultural contexts. Figure 1 provides a high-level overview of the bias landscape in LLMs considered in this review, linking upstream sources of bias to intrinsic bias, extrinsic bias in downstream tasks, cross-cutting evaluation and mitigation perspectives, and broader societal harms.
Figure 1: Overview of the bias landscape in large language models.
@@
The remainder of this paper is organized as follows. Sections 2 and 3 discuss the sources and types of intrinsic and extrinsic bias, drawing on examples from various NLP tasks to illustrate how biases arise and persist. Section 4 reviews the methodologies used to detect and measure biases, including both quantitative and qualitative approaches. Section 5 compares mitigation strategies across different stages of the model development pipeline, assessing their effectiveness and limitations. Section 6 discusses the ethical considerations and legal implications of deploying biased AI systems, underscoring the importance of fairness, transparency, and accountability in AI development. Through these discussions, we aim to contribute to the broader discourse on responsible AI by offering insights and recommendations for researchers, practitioners, and policymakers working to ensure that LLMs serve as equitable and trustworthy tools in society.
@@ 2 Intrinsic Bias
The bias presented in large language models (LLMs) can be broadly categorized into intrinsic bias and extrinsic bias based on the different stages at which the biases manifest within the model’s lifecycle and the type of bias being measured (Doan et al., 2024). Intrinsic bias refers to biases that are inherently within the internal representations or outputs of a trained or pretrained LLM and are independent of any specific downstream tasks. Extrinsic bias refers to the biases that manifest during the model’s performance on specific downstream tasks after training or fine-tuning. Intrinsic biases are typically encoded during the training or pretraining phase when the LLM learns from large-scale corpora. For instance, if a pretrained LLM consistently associates certain professions with specific genders (e.g., assuming all doctors to be male and assuming all nurses to be female) or racial stereotypes (May et al., 2019), this bias would be considered intrinsic bias. This type of bias is tied to the way the model has encoded relationships and patterns in its internal representations. In this section, we introduce the possible sources of intrinsic bias and present the manifestations of intrinsic bias.
@@
Intrinsic biases are primarily introduced during the training/pretraining stages, where models learn patterns and representations from vast corpora. These biases can be traced to several aspects of data and model design, which are outlined below.
2.1 Bias in Training Data
@@
Language models, especially LLMs that are trained on a large corpus, usually have intrinsic bias issues, since the training corpora often contain societal biases that are built into the model (Pagano et al., 2023; Ray, 2023; Goldfarb-Tarrant, 2024; Goldman and Tsotsos, 2024). These biases can arise in several ways, as described below.
@@
  • Bias from over-representativeness and under-representativeness: Certain groups, such as gender, age, race, religion, ethnicity, culture, political, and socioeconomic classes, may be under-represented or over-represented in the corpus (Das et al., 2024; Alvero et al., 2024). For example, men might be over-represented in datasets about leadership or science, while women may be more frequently mentioned in caregiving roles (UNESCO and IRCAI, 2024), thus leading to biased associations with demographic information. Another example is that the Latinx population, especially Mexican-American students, remains under-represented in U.S. higher education (Alvero et al., 2024).
@@
  • Spatial and temporal bias: LLMs trained predominantly on a corpus from certain countries or geographic locations may absorb the cultural norms and values, hence building biases into the underlying LLMs. For instance, a model trained on Western-centric data may have a skewed understanding of non-Western cultures (Ahmad and Bhattacharyya, ; Talat et al., 2022; Mukherjee et al., 2024; Lee et al., 2024; Shi et al., 2024). Prior studies have shown such effects in Arabic contexts, where models may prefer Western-associated entities over Arab ones Naous et al. (2024), and in broader multilingual settings, where performance and cultural knowledge vary substantially across under-represented languages and regions Wang et al. (2024). Similarly, data collected over different periods may reflect outdated societal norms and values. LLMs trained on these data may be biased by those outdated norms and values. For instance, historical texts may exhibit racist or sexist terminology, which the model may absorb as part of its internal representations (Salinas et al., 2023; Bussaja, 2024).
2.2 Bias in Data Collection Methods